OpenStudy (anonymous):

isolate c. a=b((1/c)-(1/d))

OpenStudy (anonymous):

so far i got (a/b)+(1/d)=(1/c)

OpenStudy (anonymous):

c= bd/ad+b

OpenStudy (anonymous):

do you mind explaining how you got that?